OpenStudy (anonymous):

PLEASE HELP: How to simplify: √98n^19

OpenStudy (anonymous):

is square root over whole thing?

OpenStudy (anonymous):

√98n^19 98 = 2 * 49 = 2 * 7 * 7 = 7√2 √n^19 = n^9√n 7n^9√2n

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

np (: